Solve for x
3x^2 = 4x-2

Answer by nitsk13(2)

3x^2 = 4x -2
3x^2 - 4x +2 = 0
X = (4 + sqr ( 16 - 4*3*2))/2*3 or X = (4 - sqr ( 16 - 4*3*2))/2*3
X = (4 + sqr ( 16 - 24))/2*3 or X = (4 - sqr ( 16 - 24))/2*3
X = (4 + sqr ( -8))/6 or X = (4 - sqr ( -8))/6
X = (4 + 2isqr2)/6 or X = (4 - 2isqr2)/6
X = (2 + isqr2)/3 or X = (2 - isqr2)/3